OpenStudy (anonymous):

Find the unit vectors that are tangent and normal to the curve at the given point (four vectors in all). Then sketch vectors and curve together. (I will put the equation and point later)

OpenStudy (anonymous):

\[y=\sum_{n=0}^{\infty}\frac{ x^n }{ n! }\] The point is (0,1)

OpenStudy (anonymous):

Do you recognize that sum as the power series representation of some nice function?

OpenStudy (anonymous):

I guess I don't. I'm pretty clueless :(

OpenStudy (anonymous):

Thats the same as e^x. Try writing out the expansion of e^x about x=0 and see

OpenStudy (anonymous):

I'm afraid I don't know how to make out the expansion either..

OpenStudy (anonymous):

The expansion of a function f about x=0 is \[\sum_{n=0}^{\infty} \frac{f^{(n)}(0)}{n!}x^{n}\] and in this case the expansion converges to the function for all x.

OpenStudy (anonymous):

so the equation equals to e^x?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

is it right when I answer the unit vector tangent of the curve is <1,1> and the unit vector normal is <1,-1>?